If sin(A+B)=1/2 and cos(A+B)=1/2....find the least values for a and b....and find the most general solution for a and b

Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Bindu,

Ans:- Let A+B=θ

Then according to the given conditions,

cos²θ+sin²θ=(1/4 + 1/4)=1/2≠1 Hence there is no solution

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.
win exciting gifts by answering the questions on Discussion Forum.

All the best Bindu !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR